When 10 is added to half the sum of a number and 8, the result is 30. what is the number?
Answer by ewatrrr(24785)


Hi
Let x represent the number
Question states***
(1/2)(n+8) + 10 = 30
Solving for x
(n+8) + 20 = 60
n = 32
CHECKING our Answer***
(1/2)40 + 10 = 30